6.) Flagstaff Enterprises expected to have free cash flow in the coming year of $8 million, and this free cash flow is expected to grow at a rate of 3% per year thereafter. Flagstaff has an equity cost of capital of 13%, a debt cost of capital of 7%, and it has a 35% corporate tax rate. If Flagstaff maintains a .5 debt to equity ratio, then Flagstaff's pre-tax WACC is closest to ________%.

  • A. 9.0
  • B. 10.0
  • C. 10.5
  • D. 11.0

Use data from Q6, if Flagstaff currently maintains a .5 debt to equity ratio, then the value of Flagstaff as an all-equity firm would be closest to $________million.

  • A. 73
  • B. 80
  • C. 100
  • D. 115

Use data from Q6, if Flagstaff currently maintains a .5 debt to equity ratio, then Flagstaff's after-tax WACC is closest to ______%.

  • A. 10.51
  • B. 10.00
  • C. 9.51
  • D. 8.75
Answer #1

SEE THE IMAGE. ANY DOUBTS, FEEL FREE TO ASK. THUMBS UP PLEASEv 1* 08-01 ENG 24-03-207027 E86 FLAGSTAFF DEBT EQUITY RATIO = 0.5 Wd = 0.5/1.5 AND We = 1/1.5 PRE TAX WACC = PRE TAX WACC =